OpenStudy (anonymous):

I NEED HELP ASAP! MEDAL, FAN! what would a cosine function look like with a given amplitude of 7, period of pi/6, no horizontal shift, and vertical shift of down 2?

OpenStudy (anonymous):

what do u need help with @countrywhitegirl1997

OpenStudy (anonymous):

answering the question above! also trhe hint for it is :f(x)= a cos(bx+c)+d

OpenStudy (anonymous):

the hint is the formula u r supost to use

OpenStudy (anonymous):

okay, well i dont know how to do it! D: I am studying and i need it kind of fast, i am doing a test with my teacher soon!

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

hank on brb

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

wow this is hard

OpenStudy (anonymous):

can you just try your best?

OpenStudy (anonymous):

im trying to watch a 3 month old as i am doing this study guide. please i need you help.

OpenStudy (misty1212):

hi

OpenStudy (misty1212):

if you are still there we can do this quickly let me know

OpenStudy (anonymous):

hey, you can help me?

OpenStudy (misty1212):

yeah sure in like 3 easy steps

OpenStudy (misty1212):

\[y=a\cos(bx)+c\] amplitude is 7 so \[y=7\cos(bx)+c\]

OpenStudy (misty1212):

vertical shift down 2 so \[y=7\cos(bx)-2\]

OpenStudy (misty1212):

period is \(\frac{\pi}{6}\) is the only part that requires algebra

OpenStudy (misty1212):

period is \(\frac{2\pi}{b}\) we need to find \(b\) so solve \[\frac{2\pi}{b}=\frac{\pi}{6}\]for \(b\)

OpenStudy (misty1212):

if you do that carefully (just do it without the \(\pi\) you should get \(b=12\) final answer \[y=7\cos(12x)-2\]

OpenStudy (misty1212):

@countrywhitegirl1997 that wasn't too bad was it?

OpenStudy (anonymous):

omg thank you so so so much! (: and no it wasnt.

OpenStudy (misty1212):

lol you are welcome :)

OpenStudy (anonymous):

@misty1212 can you help me with a few more?

OpenStudy (misty1212):

sure i got a couple minutes

OpenStudy (anonymous):

okay well the next one i need help with it: How are the special right triangle identities used to find the coordinate points around the unit circle?

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@misty1212 how do you get 12 when solving for b? i just dont get that step